# Dynamic Contextual Pricing with Doubly Non-Parametric Random Utility Models

### v1

@@ 3.2 Regret Analysis of Decision Making
The following result establishes an upper bound for the cumulative regret of Algorithm 1 over a time horizon $T$ when DNN estimator is applied in the mean utility function estimation. The proof of Theorem 7 is provided in Appendix LABEL:append:proof-regret.
@@
Assume the DNN estimator is applied with subsampling size $s=O(n_{k,exp}^{\frac{d}{d+4}})$ to estimate the mean utility function. Suppose condition of Theorem 5 for DNN estimator and Assumptions 14 hold. The regret of Algorithm 1 over time $T$ involves two settings, depending on the relationship between the dimension $d$ and the smoothness of the utility function and the market noise distribution.
  1. 1.
    When $1\leq d<\frac{6}{m-1}$ and $T\geq[\max(C_{1}d(1+\frac{\log T}{2m+1}),3\log T)]^{\frac{4m-1}{m}}$, let the exploration phase size in the $k$th episode be $n_{k,exp}=n_{k}^{\frac{2m+1}{4m-1}}$, then the total regret satisfies
    $\operatorname{Reg}(\pi,T)\leq CT^{\frac{2m+1}{4m-1}}[(\log T)^{2}+d\log T].$
  2. 2.
    When $d\geq\frac{6}{m-1}$ and $T\geq[\max(C_{1}d(1+\frac{\log T}{2m+1}),3\log T)]^{\frac{(2m+1)(d+8)}{m(d+4)}}$, let the exploration phase size in the $k$th episode be $n_{k,exp}=n_{k}^{\frac{d+4}{d+8}}$, then the total regret satisfies
    $\operatorname{Reg}(\pi,T)\leq CT^{\frac{d+4}{d+8}}[(\log T)^{2}+d\log T].$
@@
Note that the DNN estimator only requires the mean utility and density function of $\boldsymbol{x}_{t}$ are twice differentiable with bounded derivatives, while the parallel results in Theorem 8 for TDNN estimator requires fourth-order smoothness, hence achieving a lower regret bound. When the mean utility function and the Our analysis can also be extended to a broader range of nonparametric estimators for the mean utility function, such as kernel estimators, random forest, and neural network methods, provided that these estimators achieve a similar uniform convergence rate as presented in Theorem 5 for the DNN and TDNN estimators.
@@
Assume the TDNN estimator is applied with subsampling sizes $s_{1}=O(n_{k,exp}^{\frac{d}{d+8}\lor\frac{1}{7}})$ and $s_{2}\asymp s_{1}$ for mean utility function estimation. Suppose condition of Theorem 5 for TDNN estimator and Assumptions 14 hold. The regret of Algorithm 1 over time $T$ involves two settings, depending on the relationship between the dimension $d$ and the smoothness of the utility function and the market noise distribution.
  1. 1.
    When $m<1+\min(\frac{12}{d},9)$ and $T\geq[\max(C_{1}d(1+\frac{\log T}{2m+1}),3\log T)]^{\frac{4m-1}{m}}$, let the exploration phase size in the $k$th episode be $n_{k,exp}=n_{k}^{\frac{2m+1}{4m-1}}$, then the total regret satisfies
    $\operatorname{Reg}(\pi,T)\leq{C}T^{\frac{2m+1}{4m-1}}[(\log T)^{2}+d\log T].$
  2. 2.
    When $m\geq 1+\min(\frac{12}{d},9)$ and $T\geq[\max(C_{1}d(1+\frac{\log T}{2m+1}),3\log T)]^{\frac{(2m+1)}{m}\cdot\min(\frac{d+16}{d+8},\frac{13}{7})}$, let the exploration phase size in the $k$th episode be $n_{k,exp}=n_{k}^{\frac{d+8}{d+16}\lor\frac{7}{13}}$, then the total regret satisfies
    $\operatorname{Reg}(\pi,T)\leq CT^{\frac{d+8}{d+16}\lor\frac{7}{13}}[(\log T)^{2}+d\log T].$
@@ 4.1 Synthetic Data
In this subsection, we conduct large-scale simulations to illustrate the efficiency of our policy. Recall that the pricing model is characterized by (1) and (2). To elaborate, the mean utility function is set to be: $\mathring{\mu}(\boldsymbol{x}_{t})=2(x_{t1}-1)^{2}+2(x_{t2}-1)^{2}+2(x_{t3}-1)^{2}$, where the contextual covariates $\{\boldsymbol{x}_{t}\}_{t\geq 1}$ are i.i.d. bounded random vectors of dimension $d=3$, and each vector consists of independent entries uniformly distributed according to $\operatorname{Unif}(0,2)$. The noise $\{\varepsilon_{t}\}$ are i.i.d. with probability density $f(z)\propto(1/4-z^{2})\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 1/2),$ where the smoothness order $m=2$.
@@
When implementing algorithm 1, we divide the time horizon $T$ into consecutive episodes by setting the length of the $k$-th episode as $n_{k}=2^{k-1}n_{0}$ with $k\in\mathbb{N}^{+}$ and $n_{0}=200$. We further separate every episode into an exploration phase with length $|T_{k,exp}|=\lfloor(dn_{k})^{\frac{2m+1}{4m-1}}\rfloor$ induced from Theorem 7, then the exploitation phase contains the rest of the time in that episode. In the exploration phase, we sample $p_{t}$ from $\textrm{Unif}(0,B=6.5)$, since $B=6.5$ is a valid upper bound of $v_{t}$. In the exploitation phase, we set the kernels used in 11 as follows: $K(z)=(1-11/3z^{2})(1-z^{2})^{3}\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 1/2)$. In episode $k$, we set the bandwidth $b_{k}$ as $6\cdot|T_{k,exp}|^{-\frac{1}{2m+1}}$ with $m=2$ according to the assumptions in the theoretical analysis. In reality, one can also tune the bandwidth and the order of kernel $K(\cdot)$ by using cross-validation at the end of every exploration phase. Moreover, when calculating $p_{t}=\widehat{h}_{k}\circ\widehat{\mu}_{k}(\boldsymbol{x}_{t})=\widehat{\mu}_{k}(\boldsymbol{x}_{t})+\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))$, we find $\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))$ as follows: First, we look for $x\in[-1,1]$ such that $\widehat{\phi}_{k}(x)=-\widehat{\mu}_{k}(\boldsymbol{x}_{t})$ (We choose a large interval $[-1,1]$ that contains the true support of $\phi(x)$ [-0.5, 0.5], since in reality, we might only know a range of the true support). Then, we do a transformation of variable $x$ to $x(y)=-2\cdot\exp(y)/(1+\exp(y))+1$ and solve $y$ as the root of $\widehat{\phi}_{k}(x(y))+\widehat{\mu}_{k}(\boldsymbol{x}_{t})=0$ by using Newton’s method starting at $y=0$. Finally, we set $x=-2\cdot\exp(y)/(1+\exp(y))+1$ as $\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))$ and offer $p_{t}$ according to the algorithm.
@@
Through the simulation, we also compare the efficiency of our methods over two closely related approaches: the ’RMLP-2’ proposed by (Javanmard and Nazerzadeh, 2019), and the linear model based policy introduced by (Fan et al., 2022). They both address the same problem but assume the underlying mean utility function to be the linear model and, hence, use linear regression during the exploration phase. Besides, the ’RMLP-2’ additionally assumes the noise distribution belongs to a parametric function class, while given the algorithm’s lack of access to the true noise distribution, we adopt a Gaussian distribution assumption for the noise. We conduct repeated trials of Algorithm 1, the linear model based policy and the ’RMLP-2’, respectively, from $n_{1}=200$ to $n_{8}=12800$, to document the cumulative regret $\operatorname{Reg}(\pi,T)=\sum_{t=1}^{T}\operatorname{Reg}(\pi,t)$ with $T=1$ to $\sum_{k=1}^{8}2^{k-1}*200$. Consequently, the averaged regret is stored as $\text{AveReg}(T)=\operatorname{Reg}(\pi,T)/T.$ And the illustration of the distinctions in performance is detailed in several figures. Specifically, we plot $\operatorname{Reg}(\pi,T)$ against $T$ in Figure 1(a), and the averaged regret $\text{AveReg}(T)$ against $T$ in Figure 2(a). Recall from Theorem 7 that the cummulative regret $\operatorname{Reg}(\pi,T)\lesssim[(\log T)^{2}+d\log T]T^{\frac{2m+1}{4m-1}}$ when $d\leq 6/(m-1)$. Hence, we plot $\widetilde{\textrm{reg}}(T)$ against $\log(T)-\log(1500)$ in Figure 1(b), where $\widetilde{\textrm{reg}}(T):=\log(\operatorname{Reg}(T))-(2\log\log T+\log 3\log T)-(\log(\operatorname{Reg}(T=1500))-2\log\log 1500-\log 3\log 1500)$. Lastly, we plot the averaged regret of the exploitation phase against $T_{com}$ in Figure 2(b), where the Averaged regret (exploitation phase) is defined as: $\text{AveReg}({\cal T}_{com})=\sum\limits_{t\in{\cal T}_{com}}Reg(t)/T_{com}$, and $T_{com}=|{\cal T}_{com}|.$
@@
Figure 1: (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. For (a) and (b), the solid lines represent the mean $\operatorname{Reg}(\pi,T)$ and $\tilde{\text{reg}}(T)$, respectively, over 20 independent runs. The light color areas around those solid lines depict the standard error of the estimations. The blue, green, and red lines represent linear regression, TDNN, and DNN estimations in the exploration phase in Algorithm 1, respectively. And the orange line represents applying ’RMLP-2’. The dashed lines represent the benchmark whose slopes are equal to $\frac{2m+1}{4m-1}$ with $m=2$.
@@
From Figure 1, we conclude that under the specified setting, the rates of the empirical cumulative regrets produced by Algorithm 1 (as shown by the solid green and red lines) do not exceed their theoretical counterparts given in Theorems 7(as shown by the dashed lines). In many cases, the growth rates of the empirical regrets are very close to those of the theoretical line. This demonstrates the tightness of our theoretical results. We also see that the regrets we achieved are much smaller than those two approaches. As for the comparison with linear regression, TDNN and DNN estimations are robust to the mis-specification of the mean utility function, and in comparison with ’RMLP-2’, our method is robust to the mis-specification of the noise distribution since our algorithm can adapt to all noise distributions in the non-parametric class. From Figure 2, we conclude that under the same setting, Algorithm 1 incorporating TDNN estimation surpasses all other methods in performance, since as the episode length increases, the TDNN method accelerates the convergence of regret during the exploitation phase with more accurate estimation of the underline mean utility function and consequently, more accurately approaching towards noise distribution. In contrast, the decay in the averaged regret of exploitation phase for the two benchmark approaches nearly goes to zero. This occurs because the increasing size of sample points during the exploration phase does not enhance the estimation accuracy of the mean utility function due to mis-specification.

### v2

@@ 3.2 Regret Analysis of Decision Making
The following result establishes an upper bound for the cumulative regret of Algorithm 1 over a time horizon $T$ when DNN estimator is applied in the mean utility function estimation. The proofs of Theorem 8 and 9 are provided in Appendix A, respectively.
@@
Assume the DNN estimator is applied with subsampling size $s=O(n_{k,exp}^{\frac{d}{d+4}})$ to estimate the mean utility function. Suppose condition of Theorem 5 for DNN estimator and Assumptions 14 hold. Let the exploration phase size in the $k$th episode be $n_{k,exp}=n_{k}^{\frac{2m+1}{4m-1}\lor\frac{d+4}{d+8}}$ in Algorithm 1, then for any $T\geq[\max(C_{1}d(1+\frac{\log T}{2m+1}),3\log T)]^{\frac{2m+1}{m}(\frac{4m-1}{2m+1}\land\frac{d+8}{d+4})}$, the total regret satisfies
$\operatorname{Reg}(\pi,T)\leq CT^{\frac{2m+1}{4m-1}\lor\frac{d+4}{d+8}}[(\log T)^{2}+d\log T].$ (25)
@@
Theorem 8 requires that the mean utility function $\mathring{\mu}(\cdot)$ is twice differentiable and the error distribution $F(\cdot)$ has $m$-th order derivatives. The regret upper bound results from a synergistic learning of $\mathring{\mu}(\cdot)$ and $F(\cdot)$. Specifically, learning a twice-differentiable $\mathring{\mu}(\cdot)$ incurs a regret bound of $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$, while learning an $m$-th differentiable $F(\cdot)$ incurs a regret bound of $\widetilde{\cal O}(T^{\frac{d+4}{d+8}})$. Overall, the regret bound is dominated by $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$ as $1\leq d\leq\frac{6}{m-1}$ and is instead dominated by $\widetilde{{\cal O}}(T^{\frac{d+4}{d+8}})$ as $d>\frac{6}{m-1}$.
If the mean utility function $\mathring{\mu}(\cdot)$ is further fourth differentiable, the more sophisticated TDNN estimator can be utilized to estimate $\mathring{\mu}(\cdot)$ in Algorithm 1, achieving a lower regret bound due to its debiasing capability. The corresponding regret bound of incorporating the TDNN estimator are detailed in Theorem 9 below.
@@
Assume the TDNN estimator is applied with subsampling sizes $s_{1}=O(n_{k,exp}^{\frac{d}{d+8}\lor\frac{1}{7}})$ and $s_{2}\asymp s_{1}$ for mean utility function estimation. Suppose condition of Theorem 5 for TDNN estimator and Assumptions 14 hold. Let the exploration phase size in the $k$th episode be $n_{k,exp}=n_{k}^{\frac{2m+1}{4m-1}\lor\frac{d+8}{d+16}\lor\frac{7}{13}}$ in Algorithm 1, then for any $T\geq[\max(C_{1}d(1+\frac{\log T}{2m+1}),3\log T)]^{\frac{2m+1}{m}(\frac{4m-1}{2m+1}\land\frac{d+16}{d+8}\land\frac{13}{7})}$, the total regret satisfies
$\operatorname{Reg}(\pi,T)\leq CT^{\frac{2m+1}{4m-1}\lor\frac{d+8}{d+16}\lor\frac{7}{13}}[(\log T)^{2}+d\log T].$ (26)
Remark 2.
The overall regret of doubly-nonparametric learning and pricing is a trade-off between two components: the term of order $\widetilde{{\cal O}}(T^{\frac{d+8}{d+16}\lor\frac{7}{13}})$ corresponds to learning a fourth differentiable mean utility function $\mathring{\mu}(\boldsymbol{x})$, while the term of order $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$ corresponds to learning an $m$-th differentiable distribution function $F(\cdot)$.
When $F(\cdot)$ is comparatively more difficult to learn, i.e. less smooth such that $m<1+\min(\frac{12}{d},9)$, the overall regret is dominated by $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$. When $\mathring{\mu}(\boldsymbol{x})$ is comparatively more difficult to learn, i.e. higher-dimensional such that $m\geq 1+\min(\frac{12}{d},9)$, the the overall regret is dominated by $\widetilde{{\cal O}}(T^{\frac{d+8}{d+16}\lor\frac{7}{13}})$ where the term $\frac{7}{13}$ corresponds to the setting of $d=1$.
Remark 3.
Theorem 8 and 9 require only second- and fourth-order smoothness, respectively, of the true utility function $\mathring{\mu}(\cdot)$. When the smoothness of $\mathring{\mu}(\cdot)$ reaches a general level $\ell$ for an even integer $\ell$, it is possible to develop a $(\ell/2)$-scale DNN estimator analogous to the TDNN framework that achieves better convergence rates.
Remark 4.
If we only consider the explore-and-exploit algorithms and the policy set $\Pi:=\{\pi:\pi(p_{t})=\widehat{\mu}_{k}(\boldsymbol{x}_{t})+\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))\}$ with $\widehat{\phi}_{k}(z)=z-\frac{1-\widehat{F}_{k}(z)}{\widehat{F}_{k}^{(1)}(z)}$, where $\widehat{\mu}_{k}$, $\widehat{F}_{k}$, and $\widehat{F}^{(1)}_{k}$ are estimators for the mean utility function, noise distribution, and derivative of the noise distribution in the $k$-th episode, respectively, then the optimality of the regret upper bound reduces to the optimality of these estimators. Based on Stone (1982), our nonparametric estimators DNN or TDNN $\widehat{\mu}_{k}$ , Nadaraya-Watson kernel estimator $\widehat{F}$ and its derivative $\widehat{F}^{(1)}$, all achieve optimal rate of convergence. Therefore, the proposed algorithm using those estimators is minimax optimal when we only focus on the policy set $\Pi$.
Remark 5.
Our analysis can be generally extended to a broader range of nonparametric estimators for the mean utility function, such as kernel estimators, random forest, and neural network methods, provided that these estimators achieve a similar uniform convergence rate as presented in Theorem 5 for the DNN and TDNN estimators.
@@ 4.1 Synthetic Data
In this subsection, we conduct large-scale simulations to illustrate the efficiency of our policy. Recall that the pricing model is characterized by (1) and (2). To elaborate, the mean utility function is set to be $\mathring{\mu}(\boldsymbol{x}_{t})=2(x_{t1}-1)^{2}+2(x_{t2}-1)^{2}+2(x_{t3}-1)^{2}$, where the contextual covariates $\{\boldsymbol{x}_{t}\}_{t\geq 1}$ are i.i.d. bounded random vectors of dimension $d=3$, and each vector consists of independent entries uniformly distributed according to $\operatorname{Unif}(0,2)$. The noise $\{\varepsilon_{t}\}$ are i.i.d. with probability density $f(z)\propto(1/4-z^{2})\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 1/2),$ where the smoothness order $m=2$.
@@
When implementing algorithm 1, we divide the time horizon $T$ into consecutive episodes by setting the length of the $k$-th episode as $n_{k}=2^{k-1}n_{0}$ with $k\in\mathbb{N}^{+}$ and $n_{0}=200$. We further separate every episode into an exploration phase with length $|T_{k,exp}|=\lfloor(dn_{k})^{\frac{2m+1}{4m-1}}\rfloor$ induced from Theorem 8, then the exploitation phase contains the rest of the time in that episode. In the exploration phase, we sample $p_{t}$ from $\textrm{Unif}(0,B=6.5)$, since $B=6.5$ is a valid upper bound of $v_{t}$. In the exploitation phase, we set the kernels used in 12 as follows: $K(z)=(1-11/3z^{2})(1-z^{2})^{3}\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 1/2)$. In episode $k$, we set the bandwidth $b_{k}$ as $6\cdot|T_{k,exp}|^{-\frac{1}{2m+1}}$ with $m=2$ according to the assumptions in the theoretical analysis. In reality, one can also tune the bandwidth and the order of kernel $K(\cdot)$ by using cross-validation at the end of every exploration phase. Moreover, when calculating $p_{t}=\widehat{h}_{k}\circ\widehat{\mu}_{k}(\boldsymbol{x}_{t})=\widehat{\mu}_{k}(\boldsymbol{x}_{t})+\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))$, we find $\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))$ as follows: First, we look for $x\in[-1,1]$ such that $\widehat{\phi}_{k}(x)=-\widehat{\mu}_{k}(\boldsymbol{x}_{t})$ (We choose a large interval $[-1,1]$ that contains the true support of $\phi(x)$ [-0.5, 0.5], since in reality, we might only know a wider range of the true support). Then, we make a transformation of variable $x$ to $x(y)=-2\cdot\exp(y)/(1+\exp(y))+1$ and solve $y$ as the root of $\widehat{\phi}_{k}(x(y))+\widehat{\mu}_{k}(\boldsymbol{x}_{t})=0$ by using Newton’s method starting at $y=0$. Finally, we set $x=-2\cdot\exp(y)/(1+\exp(y))+1$ as $\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))$ and offer $p_{t}$ according to the algorithm.
@@
In the simulation, we also compare the efficiency of our methods with two closely related approaches: the ’RMLP-2’ proposed by (Javanmard and Nazerzadeh, 2019), and the linear model based policy introduced by (Fan et al., 2022). They both address the same problem but assume the underlying mean utility function to be the linear model and, hence, fit a linear regression during the exploration phase. Besides, the ’RMLP-2’ additionally assumes the noise distribution belongs to a parametric function class, while given the algorithm’s lack of access to the true noise distribution, we adopt a Gaussian distribution assumption for the noise. We conduct repeated trials of Algorithm 1, the linear model based policy and the ’RMLP-2’, respectively, from $n_{1}=200$ to $n_{8}=12800$, to document the cumulative regret $\operatorname{Reg}(\pi,T)=\sum_{t=1}^{T}\operatorname{Reg}(\pi,t)$ with $T=1$ to $\sum_{k=1}^{8}2^{k-1}*200$. Consequently, the averaged regret is stored as $\text{AveReg}(T)=\operatorname{Reg}(\pi,T)/T.$ And the illustration of the distinctions in performance is detailed in several figures. Specifically, we plot $\operatorname{Reg}(\pi,T)$ against $T$ in Figure 1(a), and the averaged regret $\text{AveReg}(T)$ against $T$ in Figure 2(a). Recall from Theorem 8 that the cummulative regret $\operatorname{Reg}(\pi,T)\lesssim[(\log T)^{2}+d\log T]T^{\frac{2m+1}{4m-1}}$ when $d\leq 6/(m-1)$. Hence, we plot $\widetilde{\textrm{reg}}(T)$ against $\log(T)-\log(1500)$ in Figure 1(b), where $\widetilde{\textrm{reg}}(T):=\log(\operatorname{Reg}(T))-(2\log\log T+\log 3\log T)-(\log(\operatorname{Reg}(T=1500))-2\log\log 1500-\log 3\log 1500)$. In addition, we plot the averaged regret of the exploitation phase against $T_{com}$ in Figure 2(b), where the y-axis Averaged Regret (Exploitation Phase) is defined as: $\text{AveReg}({\cal T}_{com})=\sum\limits_{t\in{\cal T}_{com}}Reg(t)/T_{com}$, and $T_{com}=|{\cal T}_{com}|.$
@@
Figure 1: (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. For (a) and (b), the solid lines represent the mean $\operatorname{Reg}(\pi,T)$ and $\tilde{\text{reg}}(T)$, respectively, over 20 independent runs. The light color areas around those solid lines depict the standard error of the estimations. The blue, green, and red lines represent linear regression, TDNN, and DNN estimations in the exploration phase in Algorithm 1, respectively. The orange line represents applying ’RMLP-2’. The dashed lines represent the benchmark whose slopes are equal to $\frac{2m+1}{4m-1}$ with $m=2$.
@@
From Figure 1, we conclude that under the specified setting, the rates of the empirical cumulative regrets produced by Algorithm 1 (as shown by the solid green and red lines) do not exceed their theoretical counterparts given in Theorems 8(as shown by the dashed lines). At majority of time points, the growth rates of the empirical regrets are very close to those of the theoretical line. This demonstrates the tightness of our theoretical results. We also see that the regrets we achieved are much smaller than those two approaches. As for the comparison with linear regression, TDNN and DNN estimations are robust to the mis-specification of the mean utility function, and in comparison with ’RMLP-2’, our method is robust to the mis-specification of the noise distribution since our algorithm can adapt to all noise distributions in the non-parametric class. From Figure 2, we conclude that under the same setting, Algorithm 1 incorporating TDNN estimation surpasses all other methods in performance, since as the episode length increases, the TDNN method achieves a lower regret during the exploitation phase with more accurate estimation of the underline mean utility function. In contrast, the averaged regret curves of exploitation phase for the two benchmark approaches exhibit minimal decay as the time horizon $T$ increases. Such phenomenon arises from model misspecification and the biased learning of the mean utility function by the two benchmark approaches.